OpenStudy (anonymous):

a new machne can make 10000 cns three times fster than an older machine. with both machines working, 10000cans be made in 9 hours. how long would it take the new machine, working alone, to make the 10000 cans

OpenStudy (anonymous):

let old machine take x hours then new takes x/3 hours working together they take 9 hours so x * (x/3) -------- = 9 x + x/3 x^2 / 3 = 9 * 4x / 3 3 x^2 = 108x x^2 = 36 x x ^2 - 36x = 0 x = 0 or 36 new ,achine takes 36 / 3 = 12 hours to make 10000 cans